# On the composable security of weak coin flipping

### v2

@@ B.1 Proof of Proposition 8
Since the protocol $\pi$ is restricted to the comb model, we can assume that $\pi_{1}$ only outputs at the set $\mathcal{T}_{\pi}^{\text{even}}=\{t_{i}\in\mathcal{T}_{\pi}\mid i\text{ is even}\}$ and $\pi_{2}$ only outputs at set $\mathcal{T}_{\pi}^{\text{odd}}=\{t_{i}\in\mathcal{T}_{\pi}\mid i\text{ is odd}\}$.
For simplicity, where the cut $\mathcal{C}$ is obvious, denote for $X=B,C,C^{\prime},C^{\prime\prime},E$
@@
where $\mathcal{T}_{\pi}$ is the poset for $\pi=(\pi_{1},\pi_{2})$ and $\mathcal{T}_{\pi^{\prime}}$ is the partial order for $\pi^{\prime}=(\pi_{2}^{\prime},\pi_{1}^{\prime})$. And denote the wire space of $X=A,D,D^{\prime},D^{\prime\prime},F$ by
@@
For $\mathcal{C}\in\mathfrak{C}(\mathcal{T}_{\pi})$, $\{|i\rangle_{\hat{A}}\}$ is a orthonormal base of space $\mathcal{F}_{A}^{\chi_{\pi}(\mathcal{C})}$, The Choi-Jamiolkowski (CJ) representation of the box $\pi_{1}$ is a sesquilinear form $R_{\pi_{1}}^{\mathcal{C}}$ satisfying
$R_{\pi_{1}}^{\mathcal{C}}(\psi_{\hat{C}}\otimes\psi_{\hat{E}}\otimes\psi_{\hat{A}},\varphi_{\hat{C}}\otimes\varphi_{\hat{E}}\otimes\varphi_{\hat{A}})=\langle\psi_{\hat{C}}|\langle\psi_{\hat{E}}|\Pi_{1}^{\mathcal{C}}(|\bar{\psi}_{\hat{A}}\rangle\langle\bar{\varphi}_{\hat{A}}|)|\varphi_{\hat{E}}\rangle|\varphi_{\hat{C}}\rangle,$ (50)
@@
For $\mathcal{C}\in\mathcal{T}_{\pi}$, assume that the wire $B,C$ are connected with respect to the bases $\{|k_{\hat{C}}\rangle\},\{|l_{\hat{C}}\rangle\}$ of $\hat{C}$ and the bases $\{|k_{\hat{B}}\rangle\},\{|l_{\hat{B}}\rangle\}$ of $\hat{B}$, the wire $D,A$ are connected with respect to the orthonormal bases $\{|p_{\hat{D}}\rangle\}$, $\{|q_{\hat{D}}\rangle\}$ of $\hat{D}$ and the orthonormal bases $\{|p_{\hat{A}}\rangle\}$, $\{|q_{\hat{A}}\rangle\}$ of $\hat{A}$. The CJ representation of $\pi_{1}\pi_{2}$ is
@@
$=\sum_{k,l,p,q}R_{\pi_{1}}^{\mathcal{C}}(k_{\hat{C}}\otimes\psi_{\hat{E}}\otimes\bar{p}_{A},l_{\hat{C}}\otimes\varphi_{\hat{E}}\otimes\bar{q}_{\hat{A}})R_{\pi_{2}}^{\mathcal{C}}(p_{\hat{D}}\otimes\psi_{\hat{F}}\otimes\bar{k}_{\hat{B}},q_{\hat{D}}\otimes\varphi_{\hat{F}}\otimes\bar{l}_{\hat{B}})$
@@
where $\bar{\alpha}_{\tilde{C}^{\prime\prime}}\in\mathcal{F}_{C^{\prime\prime}}^{\mathcal{C}^{\prime}}$ and $\bar{\alpha}_{\hat{C}^{\prime}}\in\mathcal{F}_{C^{\prime}}^{\chi(\mathcal{C}^{\prime})}$ have the same coefficients in the bases of $\{|i\rangle_{\tilde{C}^{\prime\prime}}\}$ and $\{|i\rangle_{\hat{C}^{\prime}}\}$. Since the state on space $\hat{C}^{\prime\prime}$ is always a vacuum state, we omit the arguments $\psi_{\hat{C}^{\prime\prime}},\varphi_{\hat{C}^{\prime\prime}}$ of $R_{\theta_{1}^{\mathcal{C}^{\prime}}}$ and the second term of Equation (52) in the remaining part of the paper.
For all $\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}_{\pi_{1}\theta_{1}})$, wire $C$ and $C^{\prime}$ are connected with respect the bases $\{|k_{\hat{C}^{\prime}}\rangle\}$ and $\{|l_{\hat{C}^{\prime}}\rangle\}$ such that $|k_{\hat{C}^{\prime}}\rangle\cong|k_{\hat{C}}\rangle,|l_{\hat{C}^{\prime}}\rangle\cong|l_{\hat{C}}\rangle$. the CJ representation of the composition box $\bar{\pi}_{1}=\pi_{1}\theta_{1}$ is,
@@
where $|\gamma_{\hat{C}}\rangle=\left(\sum_{k}|k_{\hat{C}}\rangle\langle k_{\tilde{C}^{\prime\prime}}|\right)|\gamma_{\tilde{C}^{\prime\prime}}\rangle$ and $|\eta_{\hat{C}}\rangle=\left(\sum_{l}|l_{\hat{C}}\rangle\langle l_{\tilde{C}^{\prime\prime}}|\right)|\eta_{\tilde{C}^{\prime\prime}}\rangle$. Equation (53) follows from Definition 18, Equation (54) follows from Equation (52) and Equation (55) follows from the sesquilinearity.
Similarly, for all $\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}_{\pi_{2}^{\prime}\theta_{2}})$, wire $D$ and $D^{\prime}$ are connected with respect the bases $\{|p_{\tilde{D}^{\prime}}\rangle\}$ and $\{|q_{\tilde{D}^{\prime}}\rangle\}$ such that $|p_{\tilde{D}^{\prime}}\rangle\cong|p_{\tilde{D}}\rangle,|q_{\tilde{D}^{\prime}}\rangle\cong|q_{\tilde{D}}\rangle$. The CJ representation of the box $\bar{\pi}^{\prime}_{2}=\pi^{\prime}_{2}\theta_{2}$ is
@@
Notice that, for all $\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}_{(\pi_{1}\|\pi^{\prime}_{2})})$, we have the relation chain of the base vectors
@@
The first column of $\cong$ follows naturally by simply substituting $t_{i}\in\mathcal{T}_{\pi}$ with $t^{\prime}_{i}\in\mathcal{T}_{\pi^{\prime}}$, The second column is defined by the connection of $\pi_{1},\pi_{2}$. The third column is defined by the connection inside $\bar{\pi}_{1}$ and $\bar{\pi}^{\prime}_{2}$. The fourth column follows from the effect of delay boxes $\theta_{1}$ and $\theta_{2}$.
Then for $\mathcal{C}=\mathcal{T}_{(\pi_{1}\|\pi_{2}^{\prime})\theta}$, we choose the bases $\{|k\rangle_{\tilde{C}^{\prime\prime}}\},\{|l\rangle_{\tilde{C}^{\prime\prime}}\},\{|k\rangle_{\hat{B}}\},\{|l\rangle_{\hat{B}}\}$ for the connection of $C^{\prime\prime},B$ and the bases $\{|p_{\hat{D}^{\prime\prime}}\rangle\},\{|q_{\hat{D}^{\prime\prime}}\rangle\},\{|p_{\tilde{A}}\rangle\},\{|q_{\tilde{A}}\rangle\}$ for the connection of $D^{\prime\prime},A$. The CJ representation of the final box $\bar{\pi}_{1}\bar{\pi^{\prime}_{2}}$ is
@@
where $\mathcal{C}_{1}^{\prime}=\mathcal{C}\cap\mathcal{T}_{\pi}$ and $\mathcal{C}_{2}^{\prime}=\mathcal{C}\cap\mathcal{T}_{\pi^{\prime}}$. Since $R_{\pi_{2}}$ and $R_{\pi^{\prime}_{2}}$ have the same effect on their respective spaces, there exist isomorphisms from $\hat{B},\hat{F},\hat{D}$ to $\tilde{B},\tilde{F},\tilde{D}$ such that
@@
Note that the isomorphisms here are exactly the ones used in the first column of Equation (59). Then we have
$R_{\bar{\pi}_{1}\bar{\pi}^{\prime}_{2}}^{\mathcal{T}_{(\pi_{1}\|\pi_{2})\theta}}(\psi_{\hat{E}}\otimes\psi_{\tilde{F}},\varphi_{\hat{E}}\otimes\varphi_{\tilde{F}})=R_{\pi_{1}\pi_{2}}^{\mathcal{T}_{\pi}}(\psi_{\hat{E}}\otimes\psi_{\hat{F}},\varphi_{\hat{E}}\otimes\varphi_{\hat{F}})$ (64)
@@ B.2 Proof of Theorem 10
We first simplify the adversarial box and prove this theorem by contradiction. Let $\alpha$ be a box connected with the parallel composition $\pi_{A}\|\eta_{A}$, as shown in Figure 8(a). For convenience, $c_{A},K_{A}$ are relabeled as $E,K$. Let $\mathcal{T}$ be the ordered set of $\pi_{A}\|\eta_{A}$. Without loss of generality, we can assume the final box $(\pi_{A}\|\eta_{A})\alpha$ is defined on the ordered set $\mathcal{T}$, because any extra element $t$ introduced by $\alpha$ has no effect on $\pi_{A}\|\eta_{A}$.
@@
Figure 8: (a) The basic connection. (b) An equivalent form of the box. The wire $X,Y,V,W$ are split into $X_{1},X_{2},Y_{1},Y_{2},V_{1},V_{2},W_{1},W_{2}$. (c) Causal order of the box $(\pi_{A}\|\eta_{A})\alpha$.
According to the condition (C1) and (C2), the wire $(V_{1},W_{1})$, $(B,D)$ and $(V_{2},W_{2})$ are effective only on $\mathcal{T}_{1}$, $\mathcal{T}_{2}$ and $\mathcal{T}_{3}$, respectively. Let $\{\Phi_{\alpha}^{\mathcal{C}}\}_{\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T})}$ be the set of CPTP maps corresponding to $\alpha$. Since we are only interested in the final output, we set $\mathcal{C}=\mathcal{T}$. By [35, Lemma 9], the map $\Phi_{\alpha}^{\mathcal{T}}$ can be decomposed into three maps $\Phi_{\alpha_{1}}^{\mathcal{T}}:V_{1}\mapsto W_{1}Q_{1}$, $\Phi_{\alpha_{2}}^{\mathcal{T}}:BQ_{1}\mapsto DQ_{2}$ and $\Phi_{\alpha_{3}}^{\mathcal{T}}:V_{2}Q_{2}\mapsto W_{2}$ that satisfy
@@
Here $\Phi_{\alpha_{1}}^{\mathcal{T}}$, $\Phi_{\alpha_{2}}^{\mathcal{T}}$ and $\Phi_{\alpha_{3}}^{\mathcal{T}}$ can be seen as the map of some boxes $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$, respectively. Similarly, the box $\eta_{A}$ is split into $\eta_{A1}$ and $\eta_{A2}$. Let $\Phi_{\eta_{A}\alpha_{1}}^{\mathcal{T}}$ and $\Phi_{\pi_{A}\alpha_{2}}^{\mathcal{T}}$ be the map of connected boxes $\eta_{A}\alpha_{1}$ and $\pi_{A}\alpha_{2}$. Denote $\bar{Q}=\mathcal{F}_{Q}^{\mathcal{T}}$, then the output of $\Phi_{\eta_{A}\alpha_{1}}^{\mathcal{T}}$ is a classical-quantum state $\rho_{\bar{Q}K}=\sum_{k}P_{K}(k)|k\rangle\langle k|\otimes\rho_{\bar{Q}}^{k}$. $\Phi_{\pi_{A}\alpha_{2}}^{\mathcal{T}}$ is a map from $\mathfrak{T}(\bar{Q})$ to the classical bit $E$, which can be represented as a binary POVM measurement $\{M_{0},M_{1}\}$. Then the conditional probability is $P_{E|K}(e|k)=\mathrm{tr}(M_{e}\rho_{\bar{Q}}^{k})$.
Assume $\exists k\in\mathbb{N},\Pr[E=1|K=k]\geq z+\epsilon$, then $\mathrm{tr}(M_{e}\rho_{\bar{Q}}^{k})\geq z+\epsilon$. Consider a box $\gamma$ that outputs state $\rho_{\bar{Q}}^{z}$ and has no input. The box $\pi_{A}\alpha_{2}\gamma$ will output $E=1$ with probability larger than $z+\epsilon$, which contradicts the stand-alone security. Therefore, the first condition in Definition 9 is satisfied. The second condition follows similarly.

### v3

@@ B.1 Proof of Proposition 8
Proposition.
Proposition 8 implies that two parties running $\pi_{1}$ and $\pi_{2}^{\prime}$ with an adversary running $\theta$ is equivalent to two parties running $\pi_{1}$ and $\pi_{2}^{\prime}$ between themselves. To achieve this, the adversary performs the man-in-the-middle attack, i.e., uses delay boxes to pass messages from $\pi_{1}$ to $\pi_{2}^{\prime}$ and pass responses from $\pi_{2}^{\prime}$ to $\pi_{1}$ subsequently. To be precise, notice that the boxes $\pi_{1}$ and $\pi_{2}^{\prime}$ work on two independent posets, i.e., the top part and bottom part in Figure 8(b), respectively. The delay boxes $\theta_{1}$ and $\theta_{2}$ are the bridges that connect the two corresponding wire spaces. Then the key point is to find proper wire space isomorphisms for the wire connections such that $\pi_{2}^{\prime}$ together with the delay boxes is a perfect emulation of $\pi_{2}$.
[…]
$t_{0}$$t^{\prime}_{0}$$t_{2}$$t^{\prime}_{2}$$t_{n}$$t^{\prime}_{n}$$t_{1}$$t^{\prime}_{1}$$t_{3}$$t^{\prime}_{3}$$\mathcal{T}_{(\pi_{1}\|\pi^{\prime}_{2})\theta}:$
Figure 8: (a) This figure shows the connected box $(\pi_{1}\|\pi_{2}^{\prime})\theta$. The wire with a blue notation indicates that the messages on this wire only stay in sets $\mathcal{T}_{\pi}^{\text{even}}$ and $\mathcal{T}_{\pi^{\prime}}^{\text{even}}$, while the red notation corresponds to the sets $\mathcal{T}_{\pi}^{\text{odd}}$ and $\mathcal{T}_{\pi^{\prime}}^{\text{odd}}$. (b) This figure explains the partition of the totally ordered set $\mathcal{T}_{(\pi_{1}\|\pi_{2}^{\prime})\theta}$ with a non-standard Hasse diagram. Specifically, The blue elements at the top form the set $\mathcal{T}_{\pi}^{\text{even}}$; The blue elements at the bottom form the set $\mathcal{T}_{\pi^{\prime}}^{\text{even}}$; The red elements at the top form the set $\mathcal{T}_{\pi}^{\text{odd}}$; The red elements at the bottom form the set $\mathcal{T}_{\pi^{\prime}}^{\text{odd}}$.
Define the sets
$\mathcal{T}_{\pi}^{\text{even}}$$\coloneqq\{t_{i}\in\mathcal{T}_{\pi}\mid i\text{ is even}\},$ $\mathcal{T}_{\pi}^{\text{odd}}$$\coloneqq\{t_{i}\in\mathcal{T}_{\pi}\mid i\text{ is odd}\},$ (55)
$\mathcal{T}_{\pi^{\prime}}^{\text{even}}$$\coloneqq\{t_{i}\in\mathcal{T}_{\pi^{\prime}}\mid i\text{ is even}\},$ $\mathcal{T}_{\pi^{\prime}}^{\text{odd}}$$\coloneqq\{t_{i}\in\mathcal{T}_{\pi^{\prime}}\mid i\text{ is odd}\}.$ (56)
Because the protocol $\pi$ is restricted to the comb model, the wires $B,C,C^{\prime},C^{\prime\prime},E$ are only effective on the set $\mathcal{T}_{\pi}^{\text{even}}$ and $\mathcal{T}_{\pi^{\prime}}^{\text{even}}$. Similarly, the wires $A,D,D^{\prime},D^{\prime\prime},F$ are only effective on the set $\mathcal{T}_{\pi}^{\text{odd}}$ and $\mathcal{T}_{\pi^{\prime}}^{\text{odd}}$. By saying wire $X$ is only effective on $\mathcal{T}^{\prime}$, we mean the state on this wire is always in the space $\mathcal{F}_{X}^{\mathcal{T}^{\prime}}$.
For convenience, where the cut $\mathcal{C}$ is obvious, denote for $X=B,C,C^{\prime},C^{\prime\prime},E$
@@
where $\mathcal{T}_{\pi}$ is the poset for $\pi=(\pi_{1},\pi_{2})$ and $\mathcal{T}_{\pi^{\prime}}$ is the partial order for $\pi^{\prime}=(\pi_{2}^{\prime},\pi_{1}^{\prime})$. Denote the wire space of $X=A,D,D^{\prime},D^{\prime\prime},F$ by
@@
For any wire $X$, we call the isomorphism between $\hat{X}$ and $\tilde{X}$, obtained by substituting $t_{i}$ with $t_{i}^{\prime}$, the natural isomorphism.
For $\mathcal{C}\in\mathfrak{C}(\mathcal{T}_{\pi})$, $\{|i\rangle_{\hat{A}}\}$ is an orthonormal base of space $\mathcal{F}_{A}^{\chi_{\pi}(\mathcal{C})}$, The Choi-Jamiołkowski (CJ) representation of the box $\pi_{1}$ is a sesquilinear form $R_{\pi_{1}}^{\mathcal{C}}$ satisfying
$R_{\pi_{1}}^{\mathcal{C}}(\psi_{\hat{C}}\otimes\psi_{\hat{E}}\otimes\psi_{\hat{A}},\varphi_{\hat{C}}\otimes\varphi_{\hat{E}}\otimes\varphi_{\hat{A}})=\langle\psi_{\hat{C}}|\langle\psi_{\hat{E}}|\Pi_{1}^{\mathcal{C}}\left(|\bar{\psi}_{\hat{A}}\rangle\langle\bar{\varphi}_{\hat{A}}|\right)|\varphi_{\hat{E}}\rangle|\varphi_{\hat{C}}\rangle,$ (61)
@@
For $\mathcal{C}\in\mathfrak{C}(\mathcal{T}_{(\pi_{1}\|\pi_{2}^{\prime})\theta})$, assume that the wire $B,C$ are connected with respect to the bases $\{|k_{\hat{C}}\rangle\},\{|l_{\hat{C}}\rangle\}$ of $\hat{C}$ and the bases $\{|k_{\hat{B}}\rangle\},\{|l_{\hat{B}}\rangle\}$ of $\hat{B}$. The wire $D,A$ are connected with respect to the orthonormal bases $\{|p_{\hat{D}}\rangle\}$, $\{|q_{\hat{D}}\rangle\}$ of $\hat{D}$ and the orthonormal bases $\{|p_{\hat{A}}\rangle\}$, $\{|q_{\hat{A}}\rangle\}$ of $\hat{A}$. These bases are predefined by the protocol $\pi$ itself. The CJ representation of $\pi_{1}\pi_{2}$ is
@@
$=\sum_{k,l,p,q}R_{\pi_{1}}^{\mathcal{C}}(k_{\hat{C}}\otimes\psi_{\hat{E}}\otimes\bar{p}_{A},l_{\hat{C}}\otimes\varphi_{\hat{E}}\otimes\bar{q}_{\hat{A}})R_{\pi_{2}}^{\mathcal{C}}(p_{\hat{D}}\otimes\psi_{\hat{F}}\otimes\bar{k}_{\hat{B}},q_{\hat{D}}\otimes\varphi_{\hat{F}}\otimes\bar{l}_{\hat{B}}).$
@@
where $\bar{\alpha}_{\tilde{C}^{\prime\prime}}\in\mathcal{F}_{C^{\prime\prime}}^{\mathcal{C}^{\prime}}$ and $\bar{\alpha}_{\hat{C}^{\prime}}\in\mathcal{F}_{C^{\prime}}^{\chi(\mathcal{C}^{\prime})}$ have the same coefficients in the bases of $\{|i\rangle_{\tilde{C}^{\prime\prime}}\}$ and $\{|i\rangle_{\hat{C}^{\prime}}\}$. Since the state in space $\hat{C}^{\prime\prime}$ is always a vacuum state, we omit the arguments $\psi_{\hat{C}^{\prime\prime}},\varphi_{\hat{C}^{\prime\prime}}$ of $R_{\theta_{1}^{\mathcal{C}^{\prime}}}$ and the second term of Equation 63 in the remaining part of the paper.
For all $\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}_{\pi_{1}\theta_{1}})$, wire $C$ and $C^{\prime}$ are connected with respect to the bases $\{|k_{\hat{C}^{\prime}}\rangle\}$ and $\{|l_{\hat{C}^{\prime}}\rangle\}$. Then we have the isomorphisms $|k_{\hat{C}^{\prime}}\rangle\cong|k_{\hat{C}}\rangle,|l_{\hat{C}^{\prime}}\rangle\cong|l_{\hat{C}}\rangle$  [^3] Since any pair of bases $\{|k_{\hat{C}^{\prime}}\rangle\}$, $\{|k_{\hat{C}}\rangle\}$ defines a specific isomorphism between $\hat{C}^{\prime}$ and $\hat{C}$, we use the notation $|k_{\hat{C}^{\prime}}\rangle\cong|k_{\hat{C}}\rangle$ to denote this isomorphism.. the CJ representation of the composition box $\bar{\pi}_{1}=\pi_{1}\theta_{1}$ is,
@@
where $|\gamma_{\hat{C}}\rangle=\left(\sum_{k}|k_{\hat{C}}\rangle\langle k_{\tilde{C}^{\prime\prime}}|\right)|\gamma_{\tilde{C}^{\prime\prime}}\rangle$ and $|\eta_{\hat{C}}\rangle=\left(\sum_{l}|l_{\hat{C}}\rangle\langle l_{\tilde{C}^{\prime\prime}}|\right)|\eta_{\tilde{C}^{\prime\prime}}\rangle$. Equation 64 follows from Definition 18, Equation 65 follows from Equation 63, and Equation 66 follows from the sesquilinearity.
Similarly, for all $\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}_{\pi_{2}^{\prime}\theta_{2}})$, wire $D$ and $D^{\prime}$ are connected with respect to bases $\{|p_{\tilde{D}^{\prime}}\rangle\}$ and $\{|q_{\tilde{D}^{\prime}}\rangle\}$. The CJ representation of the box $\bar{\pi}^{\prime}_{2}=\pi^{\prime}_{2}\theta_{2}$ is
@@
Notice that, for all $\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}_{(\pi_{1}\|\pi^{\prime}_{2})})$, we have the isomorphism chain between the wires:
@@
The isomorphisms $\cong$ in the first column are natural isomorphisms. The isomorphisms in the second column are predefined by the connection of $\pi_{1}\pi_{2}$. The isomorphisms in the third column can be chosen freely; that is, they are neither predefined nor natural. The isomorphisms in the fourth column follows from the definition of delay boxes $\theta_{1}$ and $\theta_{2}$. Therefore, the isomorphisms in the third column induce two pairs of bases for $\tilde{B},\tilde{C}^{\prime\prime}$ and two pairs of bases for $\hat{A},\hat{D}^{\prime\prime}$, which are precisely the bases we will use for $B,C^{\prime\prime}$ and $A,D^{\prime\prime}$ connections.
Then, for $\mathcal{C}=\mathcal{T}_{(\pi_{1}\|\pi_{2}^{\prime})\theta}$, we choose the bases $\{|k\rangle_{\tilde{C}^{\prime\prime}}\},\{|l\rangle_{\tilde{C}^{\prime\prime}}\},\{|k\rangle_{\hat{B}}\},\{|l\rangle_{\hat{B}}\}$ for the connection of $C^{\prime\prime},B$ and the bases $\{|p_{\hat{D}^{\prime\prime}}\rangle\},\{|q_{\hat{D}^{\prime\prime}}\rangle\},\{|p_{\tilde{A}}\rangle\},\{|q_{\tilde{A}}\rangle\}$ for the connection of $D^{\prime\prime},A$. The CJ representation of the final box $\bar{\pi}_{1}\bar{\pi}^{\prime}_{2}$ is
@@
where $\mathcal{C}_{1}^{\prime}=\mathcal{C}\cap\mathcal{T}_{\pi}$ and $\mathcal{C}_{2}^{\prime}=\mathcal{C}\cap\mathcal{T}_{\pi^{\prime}}$. $R_{\pi_{2}}$ and $R_{\pi^{\prime}_{2}}$ should have the same effect on their respective wire spaces. Rigorously speaking, for natural isomorphisms
$|p_{\tilde{D}}\rangle\cong|p_{\hat{D}}\rangle,|q_{\tilde{D}}\rangle\cong|q_{\hat{D}}\rangle,$ (74)
$|k_{\tilde{B}}\rangle\cong|k_{\hat{B}}\rangle,|l_{\tilde{B}}\rangle\cong|l_{\hat{B}}\rangle,$
$|\psi_{\tilde{F}}\rangle\cong|\psi_{\hat{F}}\rangle,|\varphi_{\tilde{F}}\rangle\cong|\varphi_{\hat{F}}\rangle,$
we have
@@
Substituting Equation 75 into Equation 73, we have
$R_{\bar{\pi}_{1}\bar{\pi}^{\prime}_{2}}^{\mathcal{T}_{(\pi_{1}\|\pi_{2})\theta}}(\psi_{\hat{E}}\otimes\psi_{\tilde{F}},\varphi_{\hat{E}}\otimes\varphi_{\tilde{F}})=R_{\pi_{1}\pi_{2}}^{\mathcal{T}_{\pi}}(\psi_{\hat{E}}\otimes\psi_{\hat{F}},\varphi_{\hat{E}}\otimes\varphi_{\hat{F}}),$ (76)
which concludes the proof.
@@ B.2 Proof of Theorem 10
Theorem.
  1. (C1)
  2. (C2)
In this proof, we focus on the first condition in Definition 9, then the second condition follows similarly.
Let $\alpha$ be the adversarial box connected to the parallel composition $\pi_{A}\|\eta_{A}$, as shown in Figure 9(a). For convenience, $c_{A},K_{A}$ in Definition 9 are relabeled as $E,K$, respectively. The idea of this proof is by contradiction, explained as follows. Our target is to bound the conditional probability $\Pr[E=1|K=k]$. Due to condition (C1) and (C2), the interaction of $\alpha$ divides into three phases (see Figure 9(b)). Observe that the only way that $K$ can influence $E$ is through the information $Q_{1}$ passed from the first phase $\alpha_{1}$ to the second phase $\alpha_{2}$. If the adversarial box $\alpha$ can bias $\Pr[E=1|K=k]$ beyond $\frac{1}{2}+\epsilon$, then the information on $Q_{1}$ conditioned on $K=k$, together with the box $\alpha_{2}$, becomes a successful attacking strategy for $\pi_{A}$ alone.
We first simplify the adversarial box and prove this theorem by contradiction. Let $\mathcal{T}$ be the ordered set of $\pi_{A}\|\eta_{A}$. Without loss of generality, we can assume the final box $(\pi_{A}\|\eta_{A})\alpha$ is defined on the ordered set $\mathcal{T}$, because any extra element $t$ introduced by $\alpha$ has no effect on $\pi_{A}\|\eta_{A}$.
@@
Figure 9: (a) The basic connection of box $(\pi_{A}\|\eta_{A})\alpha$. (b) An equivalent form of the box. The wires $X,Y,V,W$ are split into $(X_{1},X_{2}),(Y_{1},Y_{2}),(V_{1},V_{2}),(W_{1},W_{2})$, respectively. (c) Causal order of box $(\pi_{A}\|\eta_{A})\alpha$. The sets $\mathcal{T}_{1},\mathcal{T}_{2},\mathcal{T}_{3}$ as a whole are strictly ordered.
According to the condition (C1) and (C2), the wire $(V_{1},W_{1})$, $(B,D)$ and $(V_{2},W_{2})$ are effective only on $\mathcal{T}_{1}$, $\mathcal{T}_{2}$ and $\mathcal{T}_{3}$, respectively. Let $\{\Phi_{\alpha}^{\mathcal{C}}\}_{\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T})}$ be the set of CPTP maps corresponding to $\alpha$. Since we are only interested in the final output, we set $\mathcal{C}=\mathcal{T}$. By [36, Lemma 9], the map $\Phi_{\alpha}^{\mathcal{T}}$ can be decomposed into three maps $\Phi_{\alpha_{1}}^{\mathcal{T}}:V_{1}\mapsto W_{1}Q_{1}$, $\Phi_{\alpha_{2}}^{\mathcal{T}}:BQ_{1}\mapsto DQ_{2}$ and $\Phi_{\alpha_{3}}^{\mathcal{T}}:V_{2}Q_{2}\mapsto W_{2}$ that satisfy
@@
Here $\Phi_{\alpha_{1}}^{\mathcal{T}}$, $\Phi_{\alpha_{2}}^{\mathcal{T}}$ and $\Phi_{\alpha_{3}}^{\mathcal{T}}$ can be seen as the map of some boxes $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$, respectively. Similarly, the box $\eta_{A}$ is split into $\eta_{A1}$ and $\eta_{A2}$. The decomposition is illustrated in Figure 9(b).
Let $\Phi_{\eta_{A}\alpha_{1}}^{\mathcal{T}}$ and $\Phi_{\pi_{A}\alpha_{2}}^{\mathcal{T}}$ be the map of connected boxes $\eta_{A}\alpha_{1}$ and $\pi_{A}\alpha_{2}$. Denote $\bar{Q}=\mathcal{F}_{Q}^{\mathcal{T}}$, then the output of $\Phi_{\eta_{A}\alpha_{1}}^{\mathcal{T}}$ is a classical-quantum state $\rho_{\bar{Q}K}=\sum_{k}P_{K}(k)|k\rangle\langle k|\otimes\rho_{\bar{Q}}^{k}$. $\Phi_{\pi_{A}\alpha_{2}}^{\mathcal{T}}$ is a map from $\mathfrak{T}(\bar{Q})$ to the classical bit $E$, which can be represented as a binary POVM $\{M_{0},M_{1}\}$. Then the conditional probability is $P_{E|K}(e|k)=\mathrm{tr}(M_{e}\rho_{\bar{Q}}^{k})$.
Now we are ready to prove the theorem by contradiction. Assume $\exists k\in\mathbb{N},\Pr[E=1|K=k]\geq z+\epsilon$, then $\mathrm{tr}(M_{e}\rho_{\bar{Q}}^{k})\geq z+\epsilon$. Consider a box $\gamma$ that outputs state $\rho_{\bar{Q}}^{k}$ and has no input. The box $\pi_{A}\alpha_{2}\gamma$ will output $E=1$ with probability larger than $z+\epsilon$, which contradicts the stand-alone security. Therefore, the first condition in Definition 9 is satisfied.